ded Assignment I Read Chapter 3 Back to Assignment Due Wednesday 01.30.19 at 1 Keep the Highest: /17 3. Income statement Aa Aa The income statement, also known as the profit and loss (PSL) statement, provides a snapshot of the financial performance of a company during a specified period of time. It reports a firms gross income, expenses, net income, and the income that is available for distribution to its preferred and common shareholders. The income statement is prepared using the generally accepted accounting principles (GAAP) that match the firms revenues and expenses to the period which y were incurred, not necessarily when cash was received or paid. Investors and analyst wie l·infmatin gve nthe iere adem other financial statements and reports to evaluate the companys financial performance and condition. Consider the following scenario: Cold Goose Metal works Inc.s income statement reports data for its first year of operation. The finns CEO would like sales to increase by 25% neat year Cold Goose is able to achieve this level of increased sales, but its interest costs increase from 10% to 15%of earnings taxes (EBIT) 1. before nterest and 2. The companys operating costs excluding depreciation and amortization reman-70% fet sales and it depeci on and arirti ator expenses remain constant from year to year 3. The companys tax rate remains constant at 40% of its pre-tax income or earnings before taxes (EBT) 4. In Year 2, Cold Goose expects to pay $300,000 a and $768,825 of preferred and common stock dividends, respectively. Complete the then answer the questions that follow. Be sure to round each dolar value to the nearest whole 5 6 8
968,82s of pre erred and common stock dividends, respectively. the Year 2 income statement data for Cold Goose, then answer the questions that follow. Be sure to Cold GoosetMetal Works Inc. Income Statement for Year Ending December 31 Year 2 Year 1 (Forecasted) Net sales $10,000,000 Less: Operating costs, except depreciation and amortization Less: Depreciation and amortization expenses 400,000 $2,600,000 400,000 Operating income (or EBIT) Pre-tax income (or EBT) Earnings after taxes Earnings available to common shareholders Less: Interest expense 2,340,000 936,000 $1,404,000 Less: Taxes (40%) Less: Preferred stock dividends Less: Common stock dividends Contribution to retained earnings 1,104,000 631,800 $472,200 $539,675 n the results of the previous income statement calculations, complete the following statements: 5 8 9 0 K.
Earnings available to common shareholders Less: Common stock dividends Contribution to retained earnings 1,104,000 631,800 $472,200 $639,675 Given the results of the previous income statement calculations, complete the following statements In Year 2, if Cold Goose has 25,000 shares of preferred stock issued and outstanding, then each preferred share should expect to receive in annual dividends. If Cold Goose has 200,000 shares of common stock issued and outstanding, then the firms earnings per share (EPS) is expected to change · in Year 1 to in Year 2. . Cold Gooses before interest, taxes, depreciation and amortization (EBITDA) value changed from n Year 1 to in Year 2 contribution to retained earnings, $472,200 and $639,675, respectively. This is because all but one 騁atement involve payments and receipts of cash to say that Cold Gooses net inflows and outflows of cash at the end of Years 1 and 2 are equal to the companys annual of the item reported in the income 3 5
Answer #1

Answer 1:

Completed year 2 income statement is as follows:

Cold Goose Metal Works Inc. Income Statement for Year Ending December 31 Year 2 Year 1 Forecasted Net Sales Less: Operating costs, except depreciation and amortization Less: Depreciation and amortization expenses Operating income (or EBI Less: Interest expense Pre-Tax income (or EBT Less: Taxes (40% Earnings after taxes Less: Preferred stock dividends Earnings available to common stockholders Less: Common stock dividends Contribution to retained earnings $10,000,000 $12,500,000 $7,000,000 $8,750,000 $400,000 $2,600,000$3,350,000 $502,500 $2,340,000 $2,847,500 $936,000$1,139,000 $1,404,0001,708,500 $300,000 $1,104,000$1,408,500 $768,825 $639,675 $400,000 $260,000 $300,000 $631,800 $472,200

Workings:

Year 2 forecasted:

Sales = $10,000,000 * (1 + 25%) = $12,500,000

Operating costs, except depreciation and amortization = $12,500,000 * 70% = $8,750,000

Interest expense = EBIT * 15% = $3,350,000 * 15% = $502,500

Answer 2:

(A) In year 2, if Cold Goose has 25,000 shares of preferred stock issued and outstanding, then each preferred share should expect to receive $12 in annual dividends.

Workings:

Annual dividend per preference share = Total preference dividend / Preferred stock issued and outstanding

= $300,000 /25,000 = $12

(B) If Cold Goose has 200,000 shares of common stock issued and outstanding, then the firm's earnings per share (EPS) is expected to change from $3.16 in year 1 to $3.84 in year 2.

Workings:

EPS = Earnings available to common stockholders / Common stock issued and outstanding

Year 1 EPS = 631800 / 200,000 = $3.16

Year 2 EPS = 768825 / 200,000 = $3.84

(C) Cold Goose's before interest, taxes and depreciation and amortization (EBITDA) value changed from $3,000,000 year 1 to $3,750,000 in year 2.

Workings:

EBITDA = Net sales - Operating costs, except depreciation and amortization

Year 1 EBIDTA = $10,000,000 - $7,000,000 = $3,000,000

Year 2 EBIDTA = $12,500,000 - $8,750,000 = $3,750,000

(D) It is incorrect to say that Cold Goose's net inflows and outflows of cash at the end of Year 2 and Year 2 are equal to the company's annual contribution to retained earnings, $472,200 and $639,675, respectively. This is because all but one of the item reported in the income statement involve payments and receipts of cash.